find the value of 9t+6 given that -2t+1=3
Answer by checkley77(12844)

9t+6 given that -2t+1=3
-2t+1=3
-2t=3-1
-2t=2
t=2/-2
t=-1 ans.
9*-1+6
-9+6=-3 ans.